Calculate the mass of Al

produced in 1.00 hr by the

electrolysis of molten AlCl3

if

the electrical current is 10.0 A


1
Expert's answer
2021-06-28T08:11:50-0400

"Q= It"

Q= 10×60×60= 36000c

133.3g= 96500c

(36000/96500)× 133.3= 49.73g